### v3

@@ 1.3 Roadmap
In Section 2, we describe the formal problem formulation and introduce the notations that we use in the paper. In Section 3, we present the linear programming reformulation of our problem and we give a general overview of our approach to illustrate the main idea. Then, in Section 4, we describe how to construct the estimates of the LP reformulation of the CMDP problems, which will be used in our algorithm. The formal algorithm is presented in Section 5 and the corresponding theoretical analysis is conducted in Section 6.
@@ 2 Preliminaries
$V_{k}(\pi,\mu_{1})=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot c_{k}(s_{t},a_{t})\mid\mu_{1}\right]\leq\alpha_{k},\leavevmode\nobreak\ \leavevmode\nobreak\ \forall k\in[K].$ (2)
@@
We assume the existence of a stylized generative model $\mathcal{M}$. The model $\mathcal{M}$ satisfies the following condition. {assumption} For each state and action pair $(s,a)$, we can query the model $\mathcal{M}$ to obtain an observation of $r(s,a)$, $c_{k}(s,a)$ for each $k\in[K]$, and the new state $s^{\prime}\in\mathcal{S}$, where the transition from $s$ to $s^{\prime}$ follows the probability kernel $P(s^{\prime}|s,a)$ independently.
Note that in reinforcement learning for CMDP problems, querying the generative model $\mathcal{M}$ can be costly. Therefore, it is desirable to query the model $\mathcal{M}$ as less as possible, while guaranteeing the near optimality of the approximate policy. To this end, we study two measures of the performance of our policy. One is regret, i.e., conditioning on we can query the model $\mathcal{M}$ at most $N$ times, we aim to derive a policy $\pi$ and bound
$\text{Regret}_{r}(\pi,N):=V_{r}(\pi^{*},\mu_{1})-V_{r}(\pi,\mu_{1})\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }\text{Regret}_{k}(\pi,N):=V_{k}(\pi,\mu_{1})-\alpha_{k},\leavevmode\nobreak\ \forall k\in[K].$ (3)
@@
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\pi,\mu_{1})\leq\varepsilon\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }V_{k}(\pi,\mu_{1})-\alpha_{k}\leq\varepsilon,\leavevmode\nobreak\ \forall k\in[K],$ (4)
@@
Remark. Note that there are other formulations of the constraints in CMDP, for example,
$V_{k}(\pi,\mu_{1})=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot c_{k}(s_{t},a_{t})\mid\mu_{1}\right]\geq\lambda_{k},\leavevmode\nobreak\ \leavevmode\nobreak\ \forall k\in[K].$ (5)
@@ 2.1 Notations
Throughout the paper, we use $n\in[N]$ to denote the time step. For the infinite horizon discounted problem in Section 2, we can query the generative model $\mathcal{M}$ for each $(s,a)\in\mathcal{S}\times\mathcal{A}$ for one time, for a time step $n\in[N]$.
For a matrix $A$, we denote by $\mathcal{J}$ a row index set and $\mathcal{I}$ a column index set. Then, $A(:,\mathcal{I})$ denotes a sub-matrix of $A$ with columns drawn from the index set $\mathcal{I}$, and $A(\mathcal{J},:)$ denotes a sub-matrix of $A$ with rows drawn from the index set $\mathcal{J}$. Similarly, $A(\mathcal{J},\mathcal{I})$ denotes a sub-matrix of $A$ with rows drawn from $\mathcal{J}$ and columns drawn from $\mathcal{I}$. For any vector $\bm{\beta}$, we use $\bm{\beta}_{\mathcal{J}}$ to denote a sub-vector of $\bm{\beta}$ with element drawn from the index set $\mathcal{J}$.
@@ 3 Problem Reformulation and Overview of Our Approach
The infinite horizon discounted setting described in Section 2 admits a linear programming reformulation. To be specific, due to the existence of the constraints, the optimal policy of a CMDP can be randomized policies (e.g. Altman (1999)), where it is optimal to take a stochastic action given the current state. Therefore, it is convenient to represent a policy through the occupation measure, which gives us the desired linear programming reformulations of the CMDP problems.
For the infinite horizon discounted problem, the occupancy measure is defined as $q_{\pi}(s,a)$ for any state $s$, action $a$, and policy $\pi$. Note that $q_{\pi}(s,a)$ represents the total expected discounted time spent on the state-action pair $(s,a)$, under policy $\pi$, multiplied by $1-\gamma$. Then, following Altman (1999), the optimal policy (and the optimal occupancy measure) can be obtained from the following linear programming.
$V^{\mathrm{Infi}}=\leavevmode\nobreak\ \max$ $\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\hat{r}(s,a)\cdot q(s,a)$ (6a)
@@
$P(\pi^{*}(s)=a)=\left\{\begin{aligned} &\frac{q^{*}(s,a)}{\sum_{a^{\prime}\in\mathcal{A}}q^{*}(s,a^{\prime})},&&\text{if\leavevmode\nobreak\ }\sum_{a^{\prime}\in\mathcal{A}}q^{*}(s,a^{\prime})>0\\ &1/|\mathcal{A}|,&&\text{if\leavevmode\nobreak\ }\sum_{a^{\prime}\in\mathcal{A}}q^{*}(s,a^{\prime})=0,\end{aligned}\right.$ (7)
@@ 3.1 Characterization of Instance Hardness
Note that one crucial step for achieving problem-dependent bounds is to characterize the hardness of the underlying problem instance and define a gap that separates the optimal policies from the others. For multi-arm-bandit problem, the characterization of the hardness can be the gap between the optimal arm and the best sub-optimal arm (e.g. Lai and Robbins (1985)). For reinforcement learning problem, the characterization of the hardness can also be the gap between the optimal policy and the best sub-optimal policy (e.g. Auer et al. (2008)). As long as this separation gap is a positive constant, denoted by $\Delta$, separating the optimal policy from the others with a probability at least $1-\epsilon$ would require samples at most $\frac{1}{\Delta}\cdot\log(1/\epsilon)$, which finally implies the instance-optimal sample complexity bound.
@@
Since the problem hardness is characterized via corner points, it is essential to provide further characterization of the corner points. Note that in LP theory, the corner point is called basic solution and can be represented by LP basis, which involves the set of basic variables that are allowed to be non-zero, and the set of active constraints that are binding under the corresponding basic solution. For our approach, we identify one optimal basic solution and then stick to this optimal basic solution. However, there can be multiple optimal bases of the LP (8). We characterize one optimal basis and we exploit its structure during the entire horizon. We first observe the following result regarding an optimal basis to the LP (8), which follows from standard LP theory. The proof of Lemma 3.1 is relegated to Section 9 for completeness.
@@
Denote by $b$ the number of rows in the matrix $B$. Then, there exists a subset $\mathcal{J}^{*}\subset[K]$ with $K^{\prime}=|\mathcal{J}^{*}|$ and an optimal solution $\bm{q}^{*}$ to LP (8) such that there are $b+K^{\prime}$ variables in $\bm{q}^{*}$ are non-zero. Moreover, we denote by $\mathcal{I}^{*}$ the index set of the non-zero element in $\bm{q}^{*}$. Then, the optimal solution $\bm{q}^{*}$ can be uniquely determined as the solution to the linear system
$C(\mathcal{J}^{*},\mathcal{I}^{*})\bm{q}_{\mathcal{I}^{*}}=\bm{\alpha}_{\mathcal{J}^{*}},$ (9a)
$B(:,\mathcal{I}^{*})\bm{q}_{\mathcal{I}^{*}}=\bm{\mu},$ (9b)
@@ 3.2 Overview of Our Approach
In the first part, we aim to identify the set $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$, which characterizes one optimal basis of LP (6). By doing so, we also obtain the optimal actions to take given each state (and each period), by noting the relationship between the optimal actions and the optimal solution (occupancy measure) in (7). In the second part, we aim to obtain the optimal solution (occupancy measure) by solving the linear equations (9). However, since the parameters of the linear equations (9) are unknown, we cannot directly solve (9). Instead, we need to use historical samples to construct estimates of the parameters of (9) and gradually approximate the optimal solution. To this end, we seek tools from the online packing/linear programming literature (e.g. Agrawal et al. (2014), Li and Ye (2022), Vera and Banerjee (2021), Bumpensanti and Wang (2020)), where we interpret the right-hand side (RHS) of the linear equations (9) as resource capacities. Note that a crucial element in deriving the logarithmic or better problem-dependent regret in the previous literature is to let our algorithm be adaptive to the remaining resource capacities in real-time. We adopt this adaptive resolving idea in our algorithm.
@@ 4 Construct Estimates in CMDP
In this section, we describe how to construct estimates for LP (8). To this end, for a round $N_{0}$, we denote by $\mathcal{F}_{N_{0}}$ the filtration of all the information collected up to round $N_{0}$. Then, we denote by $\bar{C}_{N_{0}}$ (resp. $\bar{B}_{N_{0}}$) an estimate of the matrix $C$ (resp. $B$), constructed using the information in the set $\mathcal{F}_{N_{0}}$. We also denote by $\bar{\bm{r}}_{N_{0}}$ and estimate of $\hat{\bm{r}}$ constructed from the information in the set $\mathcal{F}_{N_{0}}$. Then, similar to Efroni et al. (2020), we can use the following LP to obtain an estimate of $V$ (8).
@@
with $\lambda_{N_{0}}$ being a parameter that we specify later. To bound the estimation gap, it is useful to bound the optimal dual solution to (8). To this end, we adopt the approach in Hiriart-Urruty and Lemaréchal (1996), Nedić and Ozdaglar (2009) that utilizes Slater’s condition, which is imposed as an assumption below. {assumption} There exists a policy $\bar{\pi}$ such that all the resource constraints are satisfied strictly. In other words, there exists an occupancy measure $\bar{\bm{q}}$ such that $B\bar{\bm{q}}=\bm{\mu}$ and $C\bar{\bm{q}}<\bm{\alpha}$. In fact, for each state $s\in\mathcal{S}$, there exists a null action that consumes no resource. The Slater point $\bar{\bm{q}}$ can be set as the policy that takes the null action given each state. The estimation error will be related to the gap between the Slater point $\bar{\bm{q}}$ and the optimal point $\bm{q}^{*}$. We then define the lower gap as
$\mathrm{Gap}_{1}(N_{0},\varepsilon)\geq V-\bar{V}_{N_{0}}$ (11)
@@
$\mathrm{Gap}_{2}(N_{0},\varepsilon)\geq\bar{V}_{N_{0}}-V$ (12)
with both inequalities (11) and (12) hold with probability at least $1-\varepsilon$.
@@ 4.1 Estimates for the Infinite Horizon Discounted Problem
$\bar{r}_{N_{0}}(s,a)=\frac{\sum_{n=1}^{N_{0}}r^{n}(s,a)}{N_{0}},\bar{c}_{k,N_{0}}(s,a)=\frac{\sum_{n=1}^{N_{0}}c^{n}_{k}(s,a)}{N_{0}},\text{\leavevmode\nobreak\ and\leavevmode\nobreak\ }\bar{P}_{N_{0}}(s^{\prime}|s,a)=\frac{\sum_{n=1}^{N_{0}}\mathbbm{1}_{s^{n}(s,a)=s^{\prime}}}{N_{0}},$ (13)
where $r^{n}(s,a)$ denotes the $n$-th observation of the reward, and $c^{n}_{k}(s,a)$ denotes the $n$-th observation of the $k$-th cost, and $s^{n}(s,a)$ denotes the $n$-th observation of the state transition for the state-action pair $(s,a)$, for $n\in[N_{0}]$. We denote by
$\mathrm{Rad}(N_{0},\varepsilon)=\sqrt{\frac{\log(2/\varepsilon)}{2N_{0}}}.$
Following the standard Hoeffding’s inequality, we know that $|\bar{r}_{N_{0}}(s,a)-\hat{r}(s,a)|$, $|\bar{c}_{k,N_{0}}(s,a)-\hat{c}_{k}(s,a)|$, and $|\bar{P}_{N_{0}}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|$ are all upper bounded by $\mathrm{Rad}(N_{0},\varepsilon)$ with probability at least $1-\varepsilon$. Then, we use the following LP with empirical estimates to approximate the value of $V^{\mathrm{Infi}}$, as given in (10). Here, $\bar{C}_{N_{0}}\in\mathbb{R}^{K\times|\mathcal{S}||\mathcal{A}|}$, and the element at the $k$-th row and $(s,a)$-th column of the matrix $\bar{C}_{N_{0}}$ is $\bar{c}_{k,N_{0}}(s,a)$. $\bar{B}_{N_{0}}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}||\mathcal{A}|}$, and the element at $s$-th row and $(s^{\prime},a^{\prime})$-th column of the matrix $\bar{B}_{N_{0}}$ is $(\delta_{s,s^{\prime}}-\gamma\cdot\bar{P}_{N_{0}}(s|s^{\prime},a^{\prime}))$. $\bar{\bm{r}}_{N_{0}}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$ with the $(s,a)$-th element being $\bar{r}_{N_{0}}(s,a)$. We have the following result, where the proof is relegated to Section 10.
@@
As long as $\lambda_{N_{0}}=\mathrm{Rad}(N_{0},\varepsilon)$, the following inequality
$V\leq\bar{V}_{N_{0}}+\mathrm{Rad}(N_{0},\varepsilon)\leq V+\mathrm{Rad}(N_{0},\varepsilon)+\frac{2\mathrm{Rad}(N_{0},\varepsilon)}{\min_{k\in[K]}\{\alpha_{k}\}}\cdot\left(1+\frac{|\mathcal{S}|}{1-\gamma}\right)+\frac{\mathrm{Rad}^{2}(N_{0},\varepsilon)}{\min_{k\in[K]}\{\alpha_{k}\}}\cdot\left(|\mathcal{S}|+\frac{|\mathcal{S}|^{2}}{1-\gamma}\right).$ (14)
holds with probability at least $1-(K|\mathcal{S}||\mathcal{A}|-|\mathcal{S}|^{2}|\mathcal{A}|)\cdot\varepsilon$.
@@
$\mathrm{Gap}_{1}(N_{0},\varepsilon)=\mathrm{Rad}(N_{0},\varepsilon)$ (15)
@@
$\mathrm{Gap}_{2}(N_{0},\varepsilon)=\frac{2\mathrm{Rad}(N_{0},\varepsilon)}{\min_{k\in[K]}\{\alpha_{k}\}}\cdot\left(1+\frac{|\mathcal{S}|}{1-\gamma}\right)+\frac{\mathrm{Rad}^{2}(N_{0},\varepsilon)}{\min_{k\in[K]}\{\alpha_{k}\}}\cdot\left(|\mathcal{S}|+\frac{|\mathcal{S}|^{2}}{1-\gamma}\right).$ (16)

### v4

@@ 1.3 Roadmap
In Section 2, we describe the formal problem formulation and introduce the notations that we use in the paper. In Section 3, we present the linear programming reformulation of our problem and we give a general overview of our approach to illustrate the main idea. Then, in Section 4, we describe how to construct the estimates of the LP reformulation of the CMDP problems, which will be used in our algorithm. The formal algorithm is presented in Section 5 and the corresponding theoretical analysis is conducted in Section 6. Our approaches are extended to the finite-horizon episodic setting, off-policy setting, and the on-policy setting in Section 7, where we obtain the similar instance-dependent guarantees.
@@ 2 Preliminaries
$V_{k}(\pi,\mu_{1})=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot c_{k}(s_{t},a_{t})\mid\mu_{1}\right]\leq\alpha_{k},\penalty 10000\ \penalty 10000\ \forall k\in[K].$ (2)
@@
We assume the existence of a stylized generative model $\mathcal{M}$. The model $\mathcal{M}$ satisfies the following condition. {assumption} For each state and action pair $(s,a)$, we can query the model $\mathcal{M}$ to obtain an observation of $r(s,a)$, $c_{k}(s,a)$ for each $k\in[K]$, and the new state $s^{\prime}\in\mathcal{S}$, where the transition from $s$ to $s^{\prime}$ follows the probability kernel $P(s^{\prime}|s,a)$ independently. In reinforcement learning for CMDP problems, querying the generative model $\mathcal{M}$ can be costly. Therefore, it is desirable to query the model $\mathcal{M}$ as less as possible, while guaranteeing the near optimality of the approximate policy. To this end, we study two measures of the performance of our policy. One is regret, i.e., conditioning on we can query the model $\mathcal{M}$ at most $N$ times, we aim to derive a policy $\pi$ and bound
$\text{Regret}_{r}(\pi,N):=V_{r}(\pi^{*},\mu_{1})-V_{r}(\pi,\mu_{1})\text{\penalty 10000\ \penalty 10000\ and\penalty 10000\ \penalty 10000\ }\text{Regret}_{k}(\pi,N):=V_{k}(\pi,\mu_{1})-\alpha_{k},\penalty 10000\ \forall k\in[K].$ (3)
@@
$V_{r}(\pi^{*},\mu_{1})-V_{r}(\pi,\mu_{1})\leq\varepsilon\text{\penalty 10000\ \penalty 10000\ and\penalty 10000\ \penalty 10000\ }V_{k}(\pi,\mu_{1})-\alpha_{k}\leq\varepsilon,\penalty 10000\ \forall k\in[K],$ (4)
@@
Note that there are other formulations of the constraints in CMDP, for example,
$V_{k}(\pi,\mu_{1})=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot c_{k}(s_{t},a_{t})\mid\mu_{1}\right]\geq\lambda_{k},\penalty 10000\ \penalty 10000\ \forall k\in[K].$ (5)
@@ 2.1 Notations
Throughout the paper, we use $n\in[N]$ to denote the time step. For the infinite horizon discounted problem in Section 2, we can query the generative model $\mathcal{M}$ for each $(s,a)\in\mathcal{S}\times\mathcal{A}$ for one time, for a time step $n\in[N]$. For a matrix $A$, we denote by $\mathcal{J}$ a row index set and $\mathcal{I}$ a column index set. Then, $A(:,\mathcal{I})$ denotes a sub-matrix of $A$ with columns drawn from the index set $\mathcal{I}$, and $A(\mathcal{J},:)$ denotes a sub-matrix of $A$ with rows drawn from the index set $\mathcal{J}$. Similarly, $A(\mathcal{J},\mathcal{I})$ denotes a sub-matrix of $A$ with rows drawn from $\mathcal{J}$ and columns drawn from $\mathcal{I}$. For any vector $\bm{\beta}$, we use $\bm{\beta}_{\mathcal{J}}$ to denote a sub-vector of $\bm{\beta}$ with element drawn from the index set $\mathcal{J}$.
@@ 3 Problem Reformulation and Overview of Our Approach
The infinite horizon discounted setting described in Section 2 admits a linear programming reformulation. To be specific, due to the existence of the constraints, the optimal policy of a CMDP can be randomized policies (e.g. altman1999constrained), where it is optimal to take a stochastic action given the current state. Therefore, it is convenient to represent a policy through the occupation measure, which gives us the desired linear programming reformulations of the CMDP problems.
For the infinite horizon discounted problem, the occupancy measure is defined as $q_{\pi}(s,a)$ for any state $s$, action $a$, and policy $\pi$. Note that $q_{\pi}(s,a)$ represents the total expected discounted time spent on the state-action pair $(s,a)$, under policy $\pi$, multiplied by $1-\gamma$. Then, following altman1999constrained, the optimal policy (and the optimal occupancy measure) can be obtained from the following linear programming.
$V^{\mathrm{Infi}}=\penalty 10000\ \max\$$\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\hat{r}(s,a)\cdot q(s,a)$ (6a)
@@
$P(\pi^{*}(s)=a)=\left\{\begin{aligned} &\frac{q^{*}(s,a)}{\sum_{a^{\prime}\in\mathcal{A}}q^{*}(s,a^{\prime})},&&\text{if\penalty 10000\ }\sum_{a^{\prime}\in\mathcal{A}}q^{*}(s,a^{\prime})>0\\ &1/|\mathcal{A}|,&&\text{if\penalty 10000\ }\sum_{a^{\prime}\in\mathcal{A}}q^{*}(s,a^{\prime})=0,\end{aligned}\right$. (7)
@@ 3.1 Characterization of Instance Hardness
Note that one crucial step for achieving problem-dependent bounds is to characterize the hardness of the underlying problem instance and define a gap that separates the optimal policies from the others. For multi-arm-bandit problem, the characterization of the hardness can be the gap between the optimal arm and the best sub-optimal arm (e.g. lai1985asymptotically). For reinforcement learning problem, the characterization of the hardness can also be the gap between the optimal policy and the best sub-optimal policy (e.g. auer2008near). As long as this separation gap is a positive constant, denoted by $\Delta$, separating the optimal policy from the others with a probability at least $1-\epsilon$ would require samples at most $\frac{1}{\Delta}\cdot\log(1/\epsilon)$, for an arbitrary $\epsilon$, which finally implies the instance-optimal sample complexity bound.
@@
Since the problem hardness is characterized via corner points, it is essential to provide further characterization of the corner points. Note that in LP theory, the corner point is called basic solution and can be represented by LP basis, which involves the set of basic variables that are allowed to be non-zero, and the set of active constraints that are binding under the corresponding basic solution. For our approach, we identify one optimal basic solution and then stick to this optimal basic solution. However, there can be multiple optimal bases of the LP (8). We characterize one optimal basis and we exploit its structure during the entire horizon. We first observe the following result regarding an optimal basis to the LP (8), which follows from standard LP theory. The proof of Lemma 3.1 is relegated to Section 10.1 for completeness.
@@
Denote by $b$ the number of rows in the matrix $B$. Then, there exists subsets $\mathcal{J}_{1}^{*}\subset[K]$, $\mathcal{J}_{2}^{*}\subset\mathcal{S}$, and a subset $\mathcal{I}^{*}\subset\mathcal{S}\times\mathcal{A}$ with $m=|\mathcal{J}_{1}^{*}|+|\mathcal{J}_{2}^{*}|=|\mathcal{I}^{*}|$ such that there exists an optimal solution $\bm{q}^{*}$ to LP (8) with the elements of $\bm{q}^{*}$ in the index set $\mathcal{I}^{*}$ are non-zero and the optimal solution $\bm{q}^{*}$ can be uniquely determined as the solution to the linear system
$C(\mathcal{J}^{*}_{1},\mathcal{I}^{*})\bm{q}_{\mathcal{I}^{*}}=\bm{\alpha}_{\mathcal{J}^{*}},$ (9a)
$B(\mathcal{J}^{*}_{2},\mathcal{I}^{*})\bm{q}_{\mathcal{I}^{*}}=\bm{\mu},$ (9b)
@@
Remark 3.2
The set $\mathcal{I}^{*}$ and $\mathcal{J}^{*}=(\mathcal{J}^{*}_{1},\mathcal{J}^{*}_{2})$ described in Lemma 3.1 may not strictly equal to the optimal basis of the LP $V$, but are enough to characterize the optimal basis. To be more precise, the set $\mathcal{I}^{*}$ actually contains all non-zero basic variables and the set $\mathcal{J}^{*}$ gives the supporting constraints for these non-zero basic variables such that their value can be uniquely determined by solving the linear equation in (9). Note that when the LP $V$ is degenerate, it can happen that one basic variable can be 0 which is not contained in $\mathcal{I}^{*}$. However, this will not influence the optimality of $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ since we only need to compute the value of the non-zero part of basic variables. Therefore, in what follows, we refer to $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ as an optimal basis of the LP $V$.
@@ 3.2 Overview of Our Approach
In the first part, we aim to identify the set $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$, which characterizes one optimal basis of LP (6). By doing so, we also obtain the optimal actions to take given each state (and each period), by noting the relationship between the optimal actions and the optimal solution (occupancy measure) in (7). In the second part, we aim to obtain the optimal solution (occupancy measure) by solving the linear equations (9). However, since the parameters of the linear equations (9) are unknown, we cannot directly solve (9). Instead, we need to use historical samples to construct estimates of the parameters of (9) and gradually approximate the optimal solution. To this end, we seek tools from the online packing/linear programming literature (e.g. agrawal2014dynamic, li2022online, vera2021bayesian, bumpensanti2020re), where we interpret the right-hand side (RHS) of the linear equations (9) as resource capacities. Note that a crucial element in deriving the logarithmic or better problem-dependent regret in the previous literature is to let our algorithm be adaptive to the remaining resource capacities in real-time. We adopt this adaptive resolving idea in our algorithm.
@@ 4 Construct Estimates in CMDP
In this section, we describe how to construct estimates for LP (8). To this end, for a round $N_{0}$, we denote by $\mathcal{F}_{N_{0}}$ the filtration of all the information collected up to round $N_{0}$. Then, we denote by $\bar{C}_{N_{0}}$ (resp. $\bar{B}_{N_{0}}$) an estimate of the matrix $C$ (resp. $B$), constructed using the information in the set $\mathcal{F}_{N_{0}}$. We also denote by $\bar{\bm{r}}_{N_{0}}$ and estimate of $\hat{\bm{r}}$ constructed from the information in the set $\mathcal{F}_{N_{0}}$. Then, similar to efroni2020exploration, we can use the following LP to obtain an estimate of $V$ (8).
@@
with $\lambda_{N_{0}}$ being a parameter that we specify later. To bound the estimation gap, it is useful to bound the optimal dual solution to (8). To this end, we adopt the approach in hiriart1996convex, nedic2009subgradient that utilizes Slater’s condition, which is imposed as an assumption below. {assumption} There exists a policy $\bar{\pi}$ such that all the resource constraints are satisfied strictly. In other words, there exists an occupancy measure $\bar{\bm{q}}$ such that $B\bar{\bm{q}}=\bm{\mu}$ and $C\bar{\bm{q}}<\bm{\alpha}$. In fact, for each state $s\in\mathcal{S}$, there exists a null action that consumes no resource. The Slater point $\bar{\bm{q}}$ can be set as the policy that takes the null action given each state. The estimation error will be related to the gap between the Slater point $\bar{\bm{q}}$ and the optimal point $\bm{q}^{*}$. For an arbitrary $\epsilon$, we then define the lower gap as
$\mathrm{Gap}_{1}(N_{0},\epsilon)\geq V-\bar{V}_{N_{0}}$ (11)
@@
$\mathrm{Gap}_{2}(N_{0},\epsilon)\geq\bar{V}_{N_{0}}-V$ (12)
with both inequalities (11) and (12) hold with probability at least $1-\epsilon$.
@@ 4.1 Estimates for the Infinite Horizon Discounted Problem
$\bar{r}_{N_{0}}(s,a)=\frac{\sum_{n=1}^{N_{0}}r^{n}(s,a)}{N_{0}},\bar{c}_{k,N_{0}}(s,a)=\frac{\sum_{n=1}^{N_{0}}c^{n}_{k}(s,a)}{N_{0}},\text{\penalty 10000\ and\penalty 10000\ }\bar{P}_{N_{0}}(s^{\prime}|s,a)=\frac{\sum_{n=1}^{N_{0}}\mathbbm{1}_{s^{n}(s,a)=s^{\prime}}}{N_{0}},$ (13)
where $r^{n}(s,a)$ denotes the $n$-th observation of the reward, and $c^{n}_{k}(s,a)$ denotes the $n$-th observation of the $k$-th cost, and $s^{n}(s,a)$ denotes the $n$-th observation of the state transition for the state-action pair $(s,a)$, for $n\in[N_{0}]$. We denote by $\mathrm{Rad}(N_{0},\epsilon)=\sqrt{\frac{\log(2/\epsilon)}{2N_{0}}}$. Following the standard Hoeffding’s inequality, we know that $|\bar{r}_{N_{0}}(s,a)-\hat{r}(s,a)|$, $|\bar{c}_{k,N_{0}}(s,a)-\hat{c}_{k}(s,a)|$, and $|\bar{P}_{N_{0}}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|$ are all upper bounded by $\mathrm{Rad}(N_{0},\epsilon)$ with probability at least $1-\epsilon$. Then, we use the following LP with empirical estimates to approximate the value of $V$, as given in (10). Here, $\bar{C}_{N_{0}}\in\mathbb{R}^{K\times|\mathcal{S}||\mathcal{A}|}$, and the element at the $k$-th row and $(s,a)$-th column of the matrix $\bar{C}_{N_{0}}$ is $\bar{c}_{k,N_{0}}(s,a)$. $\bar{B}_{N_{0}}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}||\mathcal{A}|}$, and the element at $s$-th row and $(s^{\prime},a^{\prime})$-th column of the matrix $\bar{B}_{N_{0}}$ is $(\delta_{s,s^{\prime}}-\gamma\cdot\bar{P}_{N_{0}}(s|s^{\prime},a^{\prime}))$. $\bar{\bm{r}}_{N_{0}}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$ with the $(s,a)$-th element being $\bar{r}_{N_{0}}(s,a)$. We have the following result, where the proof is relegated to Section 11.1.
@@
As long as $\lambda_{N_{0}}=\mathrm{Rad}(N_{0},\epsilon)$, the following inequality
$V\leq\bar{V}_{N_{0}}+\mathrm{Rad}(N_{0},\epsilon)\leq V+\mathrm{Rad}(N_{0},\epsilon)+\frac{2\mathrm{Rad}(N_{0},\epsilon)}{\min_{k\in[K]}\{\alpha_{k}\}}\cdot\left(1+\frac{|\mathcal{S}|}{1-\gamma}\right)+\frac{\mathrm{Rad}^{2}(N_{0},\epsilon)}{\min_{k\in[K]}\{\alpha_{k}\}}\cdot\left(|\mathcal{S}|+\frac{|\mathcal{S}|^{2}}{1-\gamma}\right).$ (14)
holds with probability at least $1-(K|\mathcal{S}||\mathcal{A}|-|\mathcal{S}|^{2}|\mathcal{A}|)\cdot\epsilon$.
@@
$\mathrm{Gap}_{1}(N_{0},\epsilon)=\mathrm{Rad}(N_{0},\epsilon)$ (15)
@@
$\mathrm{Gap}_{2}(N_{0},\epsilon)=\frac{2\mathrm{Rad}(N_{0},\epsilon)}{\min_{k\in[K]}\{\alpha_{k}\}}\cdot\left(1+\frac{|\mathcal{S}|}{1-\gamma}\right)+\frac{\mathrm{Rad}^{2}(N_{0},\epsilon)}{\min_{k\in[K]}\{\alpha_{k}\}}\cdot\left(|\mathcal{S}|+\frac{|\mathcal{S}|^{2}}{1-\gamma}\right).$ (16)